OpenStudy (anonymous):

Caleb took a 2 1/2 hour nap after soccer practice on Saturday afternoon. His younger brother slept 1 3/5 times as long as Caleb. How many hours did his brother sleep?

OpenStudy (anonymous):

2.5*1.6=4 hours

OpenStudy (anonymous):

8/5 * 5/2 = 40/10 = 4 hours